# Can Large Multimodal Models Understand Agricultural Scenes? Benchmarking with AgroMind

### v2

@@ 6 Conclusion
In this work, we present AgroMind, a comprehensive benchmark tailored for evaluating LMMs in the context of agricultural remote sensing. The benchmark is constructed from real-world data sources and comprises 27,247 question-answer pairs across 13 task types and 4 progressive cognitive levels: spatial perception, object understanding, scene understanding, and scene reasoning. All tasks are designed to reflect practical agricultural needs, enabling a systematic assessment of LMMs’ ability to interpret, infer, and reason over complex agricultural scenes. We conduct thorough evaluations of 24 representative LMMs, revealing notable performance gaps between current models and expert-level understanding in agriculture-related scenarios. AgroMind not only highlights the limitations of existing models but also provides a new foundation for advancing domain-specific multimodal intelligence. We hope this work inspires future research toward developing more capable, generalizable, and agriculture-aware large models.
@@ B.1 Data collection
Existing agricultural remote sensing benchmarks predominantly focus on single spatial resolutions, homogeneous data modalities, homogeneous QA templates, or isolated scenarios, resulting in limited scene coverage and overly simplified task designs. To bridge this gap, we systematically integrate eight public datasets with one proprietary field-plot dataset to establish a comprehensive “Sky–Air–Ground” agricultural visual-QA benchmark. Details regarding each dataset’s acquisition year, key attributes, resolution, and source are provided in Table 4. The selected datasets align with key agricultural decision pathways and span five core domains: (1) crop phenology monitoring, (2) pest and disease detection, (3) composite anomaly diagnosis via multi-sensor fusion, (4) precision plot management through field delineation, and (5) high-resolution forestry area assessment(see Fig 3 (c)). This multi-tiered coverage significantly extends the scenario breadth and task granularity of existing benchmarks in holistic agricultural assessment, thereby enabling comprehensive evaluation from plot-level operations to ecosystem-level services. By leveraging heterogeneous data sources—satellite, UAV, and ground-based sensors(multispectral/hyperspectral), we further introduce a three-dimensional evaluation framework (Spatial–Object–Scene)(see Fig 3 (b)) to rigorously assess cross-modal feature fusion and agronomic reasoning. Unlike prior work, our benchmark enables joint evaluation at the plot, farm, and regional scales and incorporates a dedicated forestry detection dimension, markedly enhancing the ecological-service assessment capabilities of agricultural RS models. The diverse, scenario-specific tasks encompassed by this benchmark (see Fig 9) lay the groundwork for robust and reproducible evaluation of large-scale multimodal models in precision agriculture.
@@
Geospatial distribution. To address spatial coverage concerns, we conducted statistical analysis of global data distribution (see Fig. 8). AgroMind intentionally spans 80 geographic regions including China, the United States, European Union nations, Brazil, Southeast Asia, and parts of Africa, reflecting agricultural diversity across crops, climates, and land-use patterns, achieving an extensive spatial coverage to date for agricultural remote sensing benchmarks.This global representation captures critical agricultural diversity across: (i) Climate regimes – from tropical monsoons (Indonesia) to Mediterranean (Italy) and continental (Ukraine); (ii) Land-use intensity – ranging from industrial monocultures (U.S. Corn Belt) to polycultures (Tanzanian farmlands); (iii) Crop phenology – covering growth cycles of many staple crops (maize, rice, oil palm, etc.). The rich geographic diversity of the dataset allows robust cross-regional generalization validation and provides foundational resources for global agricultural sustainability research.
@@
Data scene diversity. We carefully select nine datasets—including eight public datasets and one proprietary farmland plot dataset—that cover multiple core scenarios and multimodal data characteristics in agricultural remote sensing. The OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset and OilPalmUAV Datasetzheng2021growing provide high-resolution global canopy maps and oil-palm tree health assessments, respectively, allowing models to understand vegetation conditions at both macro and micro scales. In addition, the OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset encompasses ecologically diverse regions from both the Northern and Southern Hemispheres (e.g., tropical rainforests and temperate woodlands), facilitating precise quantification of forest coverage and area for large-scale forestry monitoring. Meanwhile, the Agriculture Vision ChallengeChiu_2020_CVPR and PhenoBench Datasetweyler2024pami , by providing pixel-level annotations, emphasize in-field anomaly detection and detailed crop condition analysis, enhancing model capability in handling complex scene details. To support these tasks, the Agriculture Vision ChallengeChiu_2020_CVPR combines RGB and near-infrared channels (512×512 pixels) with six binary masks (e.g., cloud shadow, double plant) to construct a multimodal anomaly diagnosis task. The PhenoBench Datasetweyler2024pami utilizes visibility maps (1024×1024 pixels) to quantify plant occlusion levels and frames tasks related to weed coexistence and plant integrity assessment. Furthermore, in the domain of fine-grained cropland management, our 7-Province Ground dataset integrates farmland vector data from multiple provinces with multispectral imagery, enabling precise field boundary delineation and dynamic coverage rate calculation. This supports tasks such as cropland area estimation and boundary recognition, significantly enhancing the practicality of models in agricultural planning and resource optimization.
@@
Multi-dimensional evaluation framework. Our benchmark employs a three-dimensional Spatial–Object–Scene framework to establish a hierarchical evaluation system for core agricultural remote-sensing tasks, systematically assessing large-scale multimodal models’ (LMMs) capabilities in localization, comprehension, and reasoning within agricultural contexts. In the Spatial dimension, both global canopy distribution from OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset and cropland vector data from the 7-Provinces Ground Dataset support regional area computation and coverage-rate statistics; the latter’s plot-boundary delineation tasks further challenge LMMs’ multi-scale geospatial parsing abilities. In the Object dimension, by leveraging instance-level annotations, the benchmark captures the characteristics of most individual agricultural targets and evaluates models’ precision in detection, classification, and counting within agricultural scenarios. Specifically, the Maize Tassel Identification employs YOLO annotations to construct 3×3 grid–based yield-prediction tasks, validating models’ dense-object localization and associative inference capabilities. Pest-classification tasks in the IP102 DatasetWu2019Insect and 2018 AI Challenge Datasets, together with plant-occlusion analysis tasks that PhenoBench Datasetweyler2024pami progressively generates from pixel-level visibility annotations, encompass an object-level evaluation spanning basic attribute discrimination to higher-order relational reasoning. In the Scene dimension, we probe models’ capacity for multi-level semantic decomposition, cross-element relational inference, and dynamic decision support in complex agricultural environments. Tasks include multi-scene coverage-rate comparison and distribution inference using OAM-TCD Dataestveitchmichaelis2024oamtcdgloballydiversedataset canopy data, as well as plant-occlusion analysis based on PhenoBench’s pixel-level maps and visibility metrics—requirements that ensure robust understanding and reasoning in interactive scenarios, thereby underpinning reliable decision support for precision agriculture.
@@
Temporal and phenological coverage. AgroMind establishes a comprehensive temporal axis essential for modeling crop phenodynamics, systematically capturing growth transitions, anomaly evolution, and seasonal patterns through multi-source heterogeneous data. Specifically: (i) Growth-stage representation leverages the OilPalmUAVzheng2021growing Dataset to construct QA pairs tracking developmental phases from growth ,maturity to aging, while competition management tasks at critical phenophases utilize PhenoBench’s visibility metricsweyler2024pami to quantify weed-crop interactions; (ii) Anomaly temporal sensitivity encodes time-critical states through spectral signatures—waterlogging detection via near-infrared absorption characteristics in Agriculture Vision ChallengeChiu_2020_CVPR —enabling cross-temporal implicit comparison; (iii) Multi-seasonal imagery combines the cultivated land status of multiple provinces in China at different time series (7 provincial ground datasets) and the Global Ecoregion Tree Monitoring Imagery (OAM-TCDveitchmichaelis2024oamtcdgloballydiversedataset ), covering the crop cycle to support phenological transition modeling. This temporal richness enables evaluation of longitudinal reasoning capabilities, particularly for yield prediction and climate adaptation analysis, establishing a foundation for multi-temporal agricultural assessment.
@@
7-Provinces Ground Dataset     2025     Multi-spectral GeoTIFF images; .shp parcel polygons; Reprojected to EPSG:3857; Coverage analysis for farmland (After further supplementation, the distribution of the dataset covers not only seven provinces in China, but also some regions in Southeast Asia and Europe.)     Variable     Internal Collection
@@ B.2 Data pre-processing details
To address the heterogeneity inherent in multi-source inputs, we propose a unified pre-processing framework characterized by both diversity and depth. The framework consists of four sequential stages—(i) multilevel annotation refinement, (ii) QA schema harmonization, (iii) logical reasoning enhancement, and (iv) task alignment—each contributing to improved data fidelity and downstream model robustness. First, given that each source employs distinct labels, formats, and pre-processing procedures, we manually refine most original annotations and a subset of images to derive secondary and tertiary annotations. For instance, pre-processing the near-infrared channel data of the Agriculture Vision ChallengeChiu_2020_CVPR collection and the semantic masks from PhenoBench Datasetweyler2024pami facilitates extraction of instance-level annotations for coverage areas and individual objects. Second, QA pair generation is managed via JSON-formatted label files: each file aggregates multiple entries, with each entry linking an image, a question, and its corresponding answer, thereby ensuring high operational efficiency. Crucially, the logical reasoning enhancement stage embeds artificially pre-defining hierarchical reasoning processes into QA design: rather than directly using raw labels as answers, we construct step-by-step reasoning paths (e.g., for canopy diameter calculation: locate bounding box → identify tree → measure pixel coverage → convert to physical scale; for weed control: quantify coverage → analyze occlusion → generate agronomic decisions). Finally, through hierarchical technology integration and task-driven optimization, the pre-processing pipeline transforms raw inputs into a sophisticated, logically coherent evaluation benchmark. This approach not only addresses the challenges of format heterogeneity but also enhances data robustness via targeted augmentation strategies, providing comprehensive support for the robust training and precise evaluation of large multimodal models.
@@
7-Provinces Ground dataset. We reprojected all GeoTIFF multispectral images and associated shapefile polygons to EPSG:3857. For each plot’s primary crop, we randomly sampled rectangular query regions at two scales (5% and 20% of the plot area) to omit areas with missing farmland. These regions were exported as PNGs and used to compute land-cover metrics—plot count, total cropland area, and coverage—to formulate cropland enumeration, area estimation, and related tasks. In a secondary selection phase, we manually curated high-fidelity cropland maps with their xmin and ymin anchored at the upper-left corner, and defined plot-boundary delineation queries by normalizing bounding boxes according to original cropland distribution labels.
@@ C.1 Human participants
Human 59.51 24.03 40.63 36.68 9.60 34.72
Random 49.41 1.68 20.96 24.67 0.50 19.25
GPT-4o openai2024gpt4o 69.65 29.88 49.72 38.67 20.58 42.98
Gemini-1.5-Flash team2023gemini 65.36 33.15 46.63 35.66 20.11 41.21
Gemini-1.5-Pro team2023gemini 70.15 26.96 45.74 37.23 11.88 38.99
Claude-3.5-Sonnet anthropic2024claude3 46.16 15.95 40.84 30.22 18.02 33.33
DeepSeek-VL2-tiny deepseekvl2 69.79 25.49 27.49 21.71 15.74 27.51
DeepSeek-VL2-small deepseekvl2 70.16 32.46 29.03 21.39 2.47 27.84
TinyLLaVA TinyLLaVA 61.30 27.05 33.70 23.40 6.75 29.88
InternVL2-2B InternVL2 61.23 26.80 30.69 20.50 11.56 28.49
InternVL2-4B InternVL2 59.60 26.32 36.88 25.15 5.34 31.46
XComposer2-4KHD XComposer2 63.52 25.46 36.22 30.90 11.06 32.31
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 50.59 27.61 36.38 20.60 9.43 31.07
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 54.06 28.12 24.45 18.75 11.84 24.61
InstructBLIP-Vicuna-7B InstrucBLIP 46.23 9.68 19.57 21.50 5.18 17.70
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 38.40 14.56 18.05 22.19 9.78 17.83
Mantis-LLaMA3-SigLIP Mantis 56.87 24.41 36.60 24.78 11.52 31.60
Mantis-Idefics2 Mantis 57.53 22.49 36.31 25.57 16.46 31.84
LLaVA-NeXT-8B LLaVA-NeXT 69.65 28.41 36.69 21.76 6.03 32.07
InternVL2-8B InternVL2 71.12 29.88 39.94 29.48 6.25 34.88
Idefics-2-8b Idefics2 65.14 28.02 34.19 20.13 11.18 30.85
LLaVA-NeXT-13B LLaVA-NeXT 60.56 27.85 36.98 20.39 5.12 31.43
InternVL2-26B InternVL2 61.52 30.15 42.48 29.85 5.81 35.84
LLaVA-NeXT-34B LLaVA-NeXT 51.40 31.02 41.90 25.78 6.87 34.93
GeoChat kuckreja2023geochat 40.30 27.70 29.73 23.40 3.75 26.15
GeoLLaVA-8K wang2025geollava 53.73 0.00 18.16 21.28 1.88 15.24

### v3

@@ 6 Conclusion
In this work, we present AgroMind, a comprehensive benchmark tailored for evaluating LMMs in the context of agricultural remote sensing. The benchmark is constructed from real-world data sources and comprises 28,482 question-answer pairs across 13 task types and 4 progressive cognitive levels: spatial perception, object understanding, scene understanding, and scene reasoning. All tasks are designed to reflect practical agricultural needs, enabling a systematic assessment of LMMs’ ability to interpret, infer, and reason over complex agricultural scenes. We conduct thorough evaluations of 24 representative LMMs, revealing notable performance gaps between current models and expert-level understanding in agriculture-related scenarios. AgroMind not only highlights the limitations of existing models but also provides a new foundation for advancing domain-specific multimodal intelligence. We hope this work inspires future research toward developing more capable, generalizable, and agriculture-aware large models.
@@ B.1 Data collection
Existing agricultural remote sensing benchmarks predominantly focus on single spatial resolutions, homogeneous data modalities, homogeneous QA templates, or isolated scenarios, resulting in limited scene coverage and overly simplified task designs. To bridge this gap, we systematically integrate nine public datasets with one proprietary field-parcel dataset to establish a comprehensive “Sky–Air–Ground” agricultural visual-QA benchmark. Details regarding each dataset’s acquisition year, key attributes, resolution, and source are provided in Table 4. The selected datasets align with key agricultural decision pathways and span five core domains: (1) crop phenology monitoring, (2) pest and disease detection, (3) composite anomaly diagnosis via multi-sensor fusion, (4) precision parcel management through field delineation, and (5) high-resolution forestry area assessment(see Fig 3 (c)). This multi-tiered coverage significantly extends the scenario breadth and task granularity of existing benchmarks in holistic agricultural assessment, thereby enabling comprehensive evaluation from parcel-level operations to ecosystem-level services. By leveraging heterogeneous data sources—satellite, UAV, and ground-based sensors(multispectral/hyperspectral), we further introduce a three-dimensional evaluation framework (Spatial–Object–Scene)(see Fig 3 (b)) to rigorously assess cross-modal feature fusion and agronomic reasoning. Unlike prior work, our benchmark enables joint evaluation at the parcel, farm, and regional scales and incorporates a dedicated forestry detection dimension, markedly enhancing the ecological-service assessment capabilities of agricultural RS models. The diverse, scenario-specific tasks encompassed by this benchmark (see Fig 9) lay the groundwork for robust and reproducible evaluation of large-scale multimodal models in precision agriculture.
@@
Geospatial distribution. To address spatial coverage concerns, we conducted statistical analysis of global data distribution (see Fig. 8). AgroMind intentionally spans 106 geographic regions including China, the United States, European Union nations, Brazil, Southeast Asia, and parts of Africa, reflecting agricultural diversity across crops, climates, and land-use patterns, achieving an extensive spatial coverage to date for agricultural remote sensing benchmarks.This global representation captures critical agricultural diversity across: (i) Climate regimes – from tropical monsoons (Indonesia) to Mediterranean (Italy) and continental (Ukraine); (ii) Land-use intensity – ranging from industrial monocultures (U.S. Corn Belt) to polycultures (Tanzanian farmlands); (iii) Crop phenology – covering growth cycles of many staple crops (maize, rice, oil palm, etc.). The rich geographic diversity of the dataset allows robust cross-regional generalization validation and provides foundational resources for global agricultural sustainability research.
@@
Data scene diversity. We carefully select ten datasets, including nine public datasets and one proprietary global parcel dataset, which cover multiple core scenarios and multimodal data characteristics in agricultural remote sensing. The OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset and OilPalmUAV Datasetzheng2021growing provide high-resolution global canopy maps and oil-palm tree health assessments, respectively, allowing models to understand vegetation conditions at both macro and micro scales. In addition, the OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset encompasses ecologically diverse regions from both the Northern and Southern Hemispheres (e.g., tropical rainforests and temperate woodlands), facilitating precise quantification of forest coverage and area for large-scale forestry monitoring. Meanwhile, the Agriculture Vision ChallengeChiu_2020_CVPR and PhenoBench Datasetweyler2024pami , by providing pixel-level annotations, emphasize in-field anomaly detection and detailed crop condition analysis, enhancing model capability in handling complex scene details. To support these tasks, the Agriculture Vision ChallengeChiu_2020_CVPR combines RGB and near-infrared channels (512×512 pixels) with six binary masks (e.g., cloud shadow, double plant) to construct a multimodal anomaly diagnosis task. The PhenoBench Datasetweyler2024pami utilizes visibility maps (1024×1024 pixels) to quantify plant occlusion levels and frames tasks related to weed coexistence and plant integrity assessment. Furthermore, in the domain of fine-grained cropland management, our 7-Province Ground dataset integrates farmland vector data from multiple provinces with multispectral imagery, enabling precise field boundary delineation and dynamic coverage rate calculation, and the CropHarvest satellite datasettseng2021cropharvest provides multispectral and hyperspectral cropland classification, supporting detailed crop type identification. This supports tasks such as cropland area estimation and boundary recognition, significantly enhancing the practicality of models in agricultural planning and resource optimization.
@@
Multi-dimensional evaluation framework. Our benchmark employs a three-dimensional Spatial–Object–Scene framework to establish a hierarchical evaluation system for core agricultural remote-sensing tasks, systematically assessing large-scale multimodal models’ (LMMs) capabilities in localization, comprehension, and reasoning within agricultural contexts. In the Spatial dimension, both global canopy distribution from OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset and cropland vector data from the Global Parcel Dataset support regional area computation and coverage-rate statistics; the latter’s parcel-boundary delineation tasks further challenge LMMs’ multi-scale geospatial parsing abilities. In the Object dimension, by leveraging instance-level annotations, the benchmark captures the characteristics of most individual agricultural targets and evaluates models’ precision in detection, classification, and counting within agricultural scenarios. Specifically, the Maize Tassel Identification employs YOLO annotations to construct 3×3 grid–based yield-prediction tasks, validating models’ dense-object localization and associative inference capabilities. Pest-classification tasks in the IP102 DatasetWu2019Insect and 2018 AI Challenge Datasets, together with plant-occlusion analysis tasks that PhenoBench Datasetweyler2024pami progressively generates from pixel-level visibility annotations, encompass an object-level evaluation spanning basic attribute discrimination to higher-order relational reasoning. In the Scene dimension, we probe models’ capacity for multi-level semantic decomposition, cross-element relational inference, and dynamic decision support in complex agricultural environments. Tasks include multi-scene coverage-rate comparison and distribution inference using OAM-TCD Dataestveitchmichaelis2024oamtcdgloballydiversedataset canopy data, as well as plant-occlusion analysis based on PhenoBench’s pixel-level maps and visibility metrics—requirements that ensure robust understanding and reasoning in interactive scenarios, thereby underpinning reliable decision support for precision agriculture.
@@
Temporal and phenological coverage. AgroMind establishes a comprehensive temporal axis essential for modeling crop phenodynamics, systematically capturing growth transitions, anomaly evolution, and seasonal patterns through multi-source heterogeneous data. Specifically: (i) Growth-stage representation leverages the OilPalmUAVzheng2021growing Dataset to construct QA pairs tracking developmental phases from growth, maturity to aging, while competition management tasks at critical phenophases utilize PhenoBench’s visibility metricsweyler2024pami to quantify weed-crop interactions; (ii) Anomaly temporal sensitivity encodes time-critical states through spectral signatures—waterlogging detection via near-infrared absorption characteristics in Agriculture Vision ChallengeChiu_2020_CVPR —enabling cross-temporal implicit comparison; (iii) Multi-seasonal imagery combines the cultivated land status of multiple provinces in China at different time series (7 provincial ground datasets), the Global Ecoregion Tree Monitoring Imagery (OAM-TCDveitchmichaelis2024oamtcdgloballydiversedataset ), and the CropHarvest datasettseng2021cropharvest with its annual satellite imaging cycle capturing spectral band evolution of land parcel types, to cover the crop cycle and support phenological transition modeling. This temporal richness enables evaluation of longitudinal reasoning capabilities, particularly for yield prediction and climate adaptation analysis, establishing a foundation for multi-temporal agricultural assessment.
@@
CropHarvest Dataset     2021     Global-scale agricultural dataset; Multi-modal remote sensing data (Sentinel-1/2, SRTM, ERA5); Binary and multiclass crop labels; Temporal sequences (12 months)     896x832, 960x896     https://zenodo.org/records/5828893
@@
Global Parcel Dataset     2025     Multi-spectral GeoTIFF images; .shp parcel polygons; Reprojected to EPSG:3857; Coverage analysis for farmland (After further supplementation, the distribution of the dataset covers not only seven provinces in China, but also some regions in Southeast Asia and Europe.)     Variable     Internal Collection
@@ B.2 Data pre-processing details
To address the heterogeneity inherent in multi-source inputs, we propose a unified pre-processing framework characterized by both diversity and depth. The framework consists of four sequential stages—(i) multilevel annotation refinement, (ii) QA schema harmonization, (iii) logical reasoning enhancement, and (iv) task alignment—each contributing to improved data fidelity and downstream model robustness. First, given that each source employs distinct labels, formats, and pre-processing procedures, we manually refine most original annotations and a subset of images to derive secondary and tertiary annotations. For instance, pre-processing the near-infrared channel data of the Agriculture Vision ChallengeChiu_2020_CVPR collection, the semantic masks from PhenoBench Datasetweyler2024pami , and the full-band satellite imagery data of the CropHarvesttseng2021cropharvest (using GDAL for reading and performing pixel-wise normalization to grayscale representation) facilitates extraction of instance-level annotations for coverage areas and individual objects. Second, QA pair generation is managed via JSON-formatted label files: each file aggregates multiple entries, with each entry linking an image, a question, and its corresponding answer, thereby ensuring high operational efficiency. Crucially, the logical reasoning enhancement stage embeds artificially pre-defining hierarchical reasoning processes into QA design: rather than directly using raw labels as answers, we construct step-by-step reasoning paths (e.g., for canopy diameter calculation: locate bounding box → identify tree → measure pixel coverage → convert to physical scale; for weed control: quantify coverage → analyze occlusion → generate agronomic decisions). Finally, through hierarchical technology integration and task-driven optimization, the pre-processing pipeline transforms raw inputs into a sophisticated, logically coherent evaluation benchmark. This approach not only addresses the challenges of format heterogeneity but also enhances data robustness via targeted augmentation strategies, providing comprehensive support for the robust training and precise evaluation of large multimodal models.
@@
CropHarvest Dataset. We ingested GEOJSON-formatted annotations containing binary and multiclass crop labels, along with spatiotemporal metadata. To enable visual question answering, we processed multi-spectral time-series imagery by decomposing each GeoTIFF into constituent spectral bands and generating normalized band mosaics through four critical steps: (i) parsing Sentinel-1/2 data using GDAL; (ii) applying per-band min-max normalization to mitigate sensor variation; (iii) constructing band mosaics while filtering degenerate cases (e.g., uniform-value bands); (iv) validating non-zero pixel ratios to exclude invalid observations— ensuring maximal geographic information retention for large-scale models. We subsequently leveraged polygon geometries and crop presence labels (is_crop) to instantiate two complementary QA modalities: binary classification via balanced queries (50% crop and 50% non-crop QA pairs) verifying agricultural land use, and spatial quantification through polygon-counting tasks assessing field fragmentation .
@@
Global Parcel Dataset. We reprojected all GeoTIFF multispectral images and associated shapefile polygons to the EPSG:3857 coordinate system for consistent spatial alignment. For cropland parcels within each study area, we generated random rectangular query regions with sizes ranging from 10% to 20% of the area’s minimum side length, ensuring strict non-overlapping using a buffer-based deduplication mechanism. These regions were exported as PNGs and used to compute land-cover metrics—parcels count, total cropland area, and coverage—to formulate cropland enumeration, area estimation, and related tasks. In a secondary selection phase, we manually curated high-fidelity cropland maps with their xmin and ymin anchored at the upper-left corner, and defined parcel-boundary delineation queries by normalizing bounding boxes according to original cropland distribution labels.
@@ C.1 Human participants
Human 53.52 23.35 38.97 40.80 5.80 33.15
Random 51.98 1.32 20.79 24.04 0.00 19.24
GPT-4o openai2024gpt4o 69.80 23.77 51.48 41.80 15.14 43.14
Gemini-1.5-Flash team2023gemini 65.35 23.96 48.45 38.25 16.33 40.92
Gemini-1.5-Pro team2023gemini 65.84 22.64 49.53 38.80 13.15 41.06
Claude-3.5-Sonnet anthropic2024claude3 65.84 25.28 41.59 32.79 18.73 37.11
DeepSeek-VL2-tiny deepseekvl2 64.36 20.94 28.40 24.59 10.76 27.51
DeepSeek-VL2-small deepseekvl2 65.84 24.72 28.80 23.22 7.57 28.08
TinyLLaVA TinyLLaVA 60.40 19.43 32.17 22.13 3.98 28.01
InternVL2-2B InternVL2 61.88 22.08 30.96 21.31 9.96 28.40
InternVL2-4B InternVL2 61.88 21.32 36.27 23.50 7.97 31.15
XComposer2-4KHD XComposer2 57.43 20.00 37.21 30.87 19.12 33.02
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 57.92 20.38 33.11 23.50 9.56 29.17
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 55.94 23.21 24.23 22.13 21.91 25.82
InstructBLIP-Vicuna-7B InstrucBLIP 37.13 6.98 18.24 24.59 7.97 17.39
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 49.01 11.32 17.50 22.40 15.14 19.01
Mantis-LLaMA3-SigLIP Mantis 56.93 19.62 37.35 26.78 4.78 31.18
Mantis-Idefics2 Mantis 53.96 16.23 36.68 24.59 12.75 30.41
LLaVA-NeXT-8B LLaVA-NeXT 59.90 22.64 36.47 22.68 11.16 31.53
InternVL2-8B InternVL2 60.89 24.15 43.88 31.69 3.98 36.30
Idefics-2-8b Idefics2 66.34 22.08 28.47 21.86 5.58 27.09
LLaVA-NeXT-13B LLaVA-NeXT 64.85 20.57 34.99 23.22 9.96 30.69
InternVL2-26B InternVL2 68.32 27.55 43.20 30.87 6.77 37.25
LLaVA-NeXT-34B LLaVA-NeXT 57.43 26.04 42.66 24.04 12.35 35.52
GeoChat kuckreja2023geochat 57.92 19.81 28.87 21.86 1.59 25.93
GeoLLaVA-8K wang2025geollava 53.47 0.00 17.09 22.40 0.80 15.73